### v4

@@ 6 Related Work
In recent years, blockchain technology has experienced significant developments due to its distinctive properties, sparking growing interest in solutions to ensure data confidentiality. Zou et al. [14] propose SPChain, a system built upon blockchain to guarantee privacy and effortless sharing of medical data. This is achieved through a proxy re-encryption scheme and local databases of medical institutions. Gan et al. [3] present a search method for encrypted medical data stored on a blockchain, resorting to an access control mechanism. Their architecture relies on: a consortium blockchain built on Ethereum for experimental analysis; several encryption methods such as probabilistic encryption and order-preserving encryption; and a non-relational database to store data. Remaining in the medical field, Miyachi and Mackey [8] introduce a framework for data privacy based on both on-chain and off-chain systems. They leverage consortium blockchains and asymmetric cryptography, while storing data off-chain. Unlike [3, 8, 14], MARTSIA operates on public blockchains to store IPFS locators; encrypts with MA-ABE; and stores data on IPFS.
@@ 7.0.1 Acknowledgements
This work was partly funded by projects SERICS (PE00000014) under the NRRP MUR program funded by the EU-NGEU, PINPOINT (B87G22000450001) under the PRIN MUR program, and Healt-e-Data, funded by the EU-NGEU under the Cyber 4.0 NRRP MIMIT programme.

### v5

@@ 6 Related Work
In recent years, blockchain technology has experienced significant developments due to its distinctive properties, sparking growing interest in solutions to ensure data confidentiality. Zou et al. [14] propose SPChain, a system built upon blockchain to guarantee privacy and effortless sharing of medical data. This is achieved through a proxy re-encryption scheme and local databases of medical institutions. Gan et al. [3] present a search method for encrypted medical data stored on a blockchain, resorting to an access control mechanism. Their architecture relies on: a consortium blockchain built on Ethereum for experimental analysis; several encryption methods such as probabilistic encryption and order-preserving encryption; and a non-relational database to store data. Remaining in the medical field, Miyachi and Mackey [8] introduce a framework for data privacy based on both on-chain and off-chain systems. They leverage consortium blockchains and asymmetric cryptography, while storing data off-chain. Unlike [3, 8, 14], MARTSIA operates on public blockchains to store IPFS locators, encrypts sensitive information with MA-ABE, and stores data on IPFS.
@@ 7.0.1 Acknowledgements
This work was partly funded by projects SERICS (PE00000014) under the NRRP MUR program funded by the EU-NGEU, PINPOINT (B87G22000450001) under the PRIN MUR program, and Health-e-Data, funded by the EU-NGEU under the Cyber 4.0 NRRP MIMIT programme.